Employees of a discount appliance store receive an additional 20% off of the lowest price on an item. If an employee purchases a dishwasher during a 15% off sale, how much will he pay if the dishwasher originally cost $450?
Answer by richwmiller(17219)

1-.20=.80
1-.15=.85
.80*.85*450=$306 discounted price
450-306=$144 discount
another way to do it
.15*450=$67.50
450.00-67.50=$382.50
.2*382.50=$76.50
382.50-76.50=$306 discounted price
67.50+76.50=$144 discount